# A Perspective for Adapting Generalist AI to Specialized Medical AI Applications and Their Challenges

### v1

@@ 3 Use cases of LLMs for medical AI applications
Table 1 lists a a number of studies that have explored adapting LLMs to medical AI tasks. In this section, we review a few example use cases, reinterpreting them through the lens of the proposed framework. Additionally, we discuss some potential improvements of those use cases.
@@
  • EMR integration: We can integrate LLMs with programming tools, such as an SQL interface to the underlying EMR database. This enhancement allows LLMs to dynamically access and review previous notes, only extracting necessary pieces to support the note generation process. In addition, this capability also enable the development of provider copilot to assist the retrieval of necessary patient information at their finger tips in a natural language manner, and empower them to focus more on care rather than system navigation.
@@
  • Task assistance: AI agent can further assist providers to accomplish the specific tasks and reduce their administration workload. For example, follow-up emails is important to reinforce care plans, clarify medical instructions, and patient education, etc. LLM agent can create a follow-up email that includes a summary of the visit, test results, and next steps to ensure that patients understand their health information, adhere to treatment plan, and feel supported in their care journey.
@@ 4 Opportunities and challenges in LLMs for medical AI
Multimodality
@@
Multimodal capabilities represent a key growth direction for LLMs in medical applications. A patient journey naturally comprises many information-rich modalities such as lab tests, imaging, genomics, etc. While generalist AI has demonstrated amazing capabilities in understanding and reasoning with biomedical text (e.g., MedPrompt [95]), competence gaps abound in the multimodal space, with vision-language model being a prominent example. E.g., GPT-4V performs poorly on identifying key findings from chest X-rays, even compared to much smaller domain-specific models (e.g., LLaVA-Rad [96]). Efficiently bridging such multimodal competency gaps thus represents a key growth frontier for medical AI. Progress is particularly fast in biomedical imaging, from harnessing public image-text data (e.g., BiomedCLIP [97], PLIP [98], CONCH [99]) to efficiently training vision-language models (e.g., LLaVA-Med [100]) to learn text-guided image generation and segmentation (e.g., BiomedJourney [101], RoentGen [102], BiomedParse). While promising, challenges abound in multimodal medical AI, from pretraining in challenging modalities (e.g., GigaPath [103]) to multimodal reasoning for precision health. Much remains to be done.
Interpretability
While LLMs hold significant potential for enhancing medical tasks, their interpretability remains a critical challenge. As deep learning models, LLMs often function as black boxes, making it difficult to trace the reasoning behind specific outputs. This lack of transparency can hinder clinical decision-making and raise concerns about accountability. To address these issues, developing explainable AI methodologies is essential for gaining a deeper understanding of LLMs [104]. At present, techniques such as chain-of-thought and program-of-thought can be employed to reveal the step-by-step reasoning and operational processes behind LLM outputs, improving their interpretability in medical applications [28, 105].
Hallucination
The risk of “hallucinations” in LLM outputs, where the model generates plausible-sounding but incorrect or fabricated information, has been widely documented [106]. In medicine, such errors can have serious consequences, including incorrect diagnoses, inappropriate treatments, and flawed research conclusions [61]. To mitigate hallucinations, strategies such as leveraging RAG to enforce the grounding citation, finetuning on high-quality, domain-specific data, implementing robust validation checks, and incorporating human-in-the-loop processes where medical professionals review and correct AI-generated content are essential. Additionally, quantifying the uncertainties in LLM outputs can serve as an indicator for detecting hallucinations during the application phase [107].
Regulation and compliance
The development, deployment and production of LLMs in medical practice must adhere to strict regulatory standards and compliance requirements to ensure patient safety and data privacy [108]. Regulatory bodies such as the FDA, EMA, and others have established guidelines for medical devices, which increasingly include AI-driven tools [109]. Developers of medical AI applications must navigate these regulations, establish robust practices, and process procedures to ensure that their models are validated, transparent, and compliant with relevant laws, such as HIPAA in the U.S. Additionally, as AI applications are continuously updated and retrained, maintaining compliance over time requires ongoing monitoring, documentation, and potentially re-certification.
Security and privacy
@@
The development and deployment of medical AI applications involve multiple stakeholders, including medical data providers (such as health systems, pharmaceutical companies, and real-world data companies), model providers (like healthcare AI startups), and users (comprising clinicians, patients, and these data-providing companies). AI developers play a crucial role in constructing pipelines that process data from these data providers into actionable knowledge, which is then made accessible to users by leveraging intelligence from model providers. It has been noted that LLMs are vulnerable to attacks using malicious prompts that aim to extract their training data [110]. When fine-tuned medical LLMs are used, one way to mitigate this risk is by incorporating synthetic training data [111]. Moreover, when AI developers handle data from multiple providers and present it to various users, it is essential to implement access controls within the AI applications. These controls ensure that users can only access specific data sets relevant to their queries. In addition, protecting user inputs is crucial, especially if they are utilized to optimize the pipeline and stored in prompts or databases for RAG. LLMs must prevent inadvertently disclosing one user’s data to another, maintaining strict confidentiality in user interactions.
Validation and monitoring
@@
A medical AI application comprises various modules and pipelines, each of which has the potential to malfunction, posing challenges in output assessment and debugging. Human oversight can be integrated to enhance validation, allowing users to confirm the accuracy of outputs through the provided references [21]. Furthermore, validation processes can be automated by leveraging LLM capabilities. For example, assertions can be embedded within the pipeline, enabling LLMs to self-correct their outputs during inference [22]. The increasing complexity of LLM-based systems, consisting of multiple interconnected components, can overwhelm any individual’s capacity to manage the entire architecture. LangSmith [112] is a DevOps tool for AI applications that aims to support deploying and monitoring LLM-powered applications. It allows developers to visualize traces and debug problematic components. Additionally, it facilitates the collection and extraction of erroneous inputs and outputs, which can then be utilized to build and augment validation datasets to enhance the applications.
Maintenance
Maintaining medical AI applications entails challenges to ensure the system’s stability and reliability. First, the variability of the underlying LLMs is a critical concern. Each version update can alter these models’ behavior and capabilities, potentially disrupting the system’s functionality [113]. Second, one component in the application can affect the overall system performance and behavior. This complex interconnectivity requires rigorous testing and calibration whenever any modifications are made. Last, the system must manage distribution shifts, particularly when handling edge cases. One effective strategy for addressing maintenance challenges involves constructing development datasets that allow for the evaluation and finetuning of AI pipelines. These datasets can be derived from real-world tasks or simulated scenarios enabled by LLMs [114]. Currently, creating such datasets typically depends on heuristic methods, highlighting a need for more research.
Continual optimization
@@ 5 Conclusion
Generalist AI models have the potential to revolutionize the medical field. A range of adaptation methods have been proposed to tailor these models for specialized applications (Table 1). In this review, we documented existing adaptation strategies and organized them within a framework designed to optimize the performance of medical AI applications from a systems engineering perspective. Our analysis and discussion of published use cases demonstrate the benefits of this framework as a systematic approach to developing and optimizing LLM-based medical AI. However, we also recognize the potential challenges that arise as the complexity of medical AI applications increases, particularly in monitoring, validation, and maintenance. Future research and development are essential to solidify the utility of LLM-driven medical AI applications, enhance patient outcomes, democratize access to quality healthcare, and reduce the workload on medical professionals.

### v3

@@ 3 Use cases of LLMs for medical AI applications
Table 1 lists a number of studies that have explored adapting LLMs to medical AI tasks. In this section, we review a few example use cases, reinterpreting them through the lens of the proposed framework. Additionally, we discuss some potential improvements of those use cases.
@@
  • EMR integration: We can integrate LLMs with programming tools, such as an SQL interface to the underlying EMR database. This enhancement allows LLMs to dynamically access and review previous notes, only extracting necessary pieces to support the note generation process. In addition, this capability also enables the development of provider copilots to assist in the retrieval of necessary patient information at their fingertips in a natural language manner and empower them to focus more on care rather than system navigation.
@@
  • Task assistance: AI agents can further assist providers in accomplishing specific tasks and reduce their administration workload. For example, follow-up emails are important to reinforce care plans, clarify medical instructions, and patient education, etc. LLM agents can create a follow-up email that includes a summary of the visit, test results, and next steps to ensure that patients understand their health information, adhere to treatment plans, and feel supported in their care journey.
@@ 4 Opportunities and challenges in LLMs for medical AI
Here, we describe the challenges that need to be addressed to embody the benefits of adapting LLMs to medical AI, with a focus on three critical areas: enhancing multimodal capabilities, ensuring trustworthiness and compliance, and managing system lifecycle through evaluation and continuous optimization.
4.1 Multimodality
@@
Multimodal capabilities represent a key growth direction for LLMs in medical applications. A patient journey naturally comprises many information-rich modalities such as lab tests, imaging, genomics, etc. While generalist AI has demonstrated amazing capabilities in understanding and reasoning with biomedical text (e.g., MedPrompt [100]), competence gaps abound in the multimodal space, with vision-language model being a prominent example. E.g., GPT-4V performs poorly on identifying key findings from chest X-rays, even compared to much smaller domain-specific models (e.g., LLaVA-Rad [101]). Efficiently bridging such multimodal competency gaps thus represents a key growth frontier for medical AI. Progress is particularly fast in biomedical imaging, from harnessing public image-text data (e.g., BiomedCLIP [102], PLIP [103], CONCH [104]) to efficiently training vision-language models (e.g., LLaVA-Med [105]) to learn text-guided image generation and segmentation (e.g., BiomedJourney [106], RoentGen [107], BiomedParse). While promising, challenges abound in multimodal medical AI, from pretraining in challenging modalities (e.g., GigaPath [108]) to multimodal reasoning for precision health.
4.2 Trustworthiness and compliance
While adapting powerful AI models like LLMs to medicine holds great promise, significant challenges persist in building trustworthiness and ensuring compliance with their use. Here, we identified several challenges: (1) hallucinations, (2) privacy risks, (3) explainability, and (4) regulations.
The risk of “hallucinations” in LLM outputs, where the model generates plausible but incorrect or fabricated information, is widely mentioned [109]. In the medical domain, such errors can have severe consequences, including misdiagnoses, inappropriate treatments, and flawed research conclusions [110]. To address hallucinations, retrieval-augmented generation (RAG) is often employed to guide LLM outputs and ground them in verifiable citations [111]. However, failure modes persist, such as LLMs citing incorrect sources or hallucinating citations altogether [112]. Future efforts should focus on fine-tuning models with high-quality, domain-specific datasets, implementing rigorous validation mechanisms, and incorporating human-in-the-loop workflows [113], where medical professionals review and correct AI-generated content to ensure accuracy and reliability.
@@ 4.2 Trustworthiness and compliance
The development and deployment of medical AI applications involve multiple stakeholders, including medical data providers (such as health systems, pharmaceutical companies, and real-world data companies), model providers (like healthcare AI startups), and users (comprising clinicians, patients, and these data-providing companies). AI developers play a crucial role in constructing pipelines that process data from these data providers into actionable knowledge, which is then made accessible to users by leveraging intelligence from model providers. It has been noted that LLMs are vulnerable to attacks using malicious prompts that aim to extract their training data [114]. This risk necessitates the deidentification of the PHI information from training data. When now real data can be acquired, synthetic training data can be incorporated [115, 116], which can be further protected with differential privacy [117]. Moreover, when AI developers handle data from multiple providers and present it to various users, it is essential to implement access controls within the AI applications. These controls ensure that users can only access specific data sets relevant to their queries. In addition, protecting user inputs is crucial, especially if they are utilized to optimize the pipeline and stored in prompts or databases for RAG. LLMs must prevent inadvertently disclosing one user’s data to another, maintaining strict confidentiality in user interactions.
LLM’s interpretability remains a critical challenge. It appears to show “emergent” intelligence but that property may also disappear when we take different evaluation metrics [118]. As deep learning models, LLMs often function as black boxes, making it difficult to trace the reasoning behind specific outputs. This lack of transparency can hinder clinical decision-making and raise concerns about accountability. To address these issues, developing explainable AI methodologies is essential for gaining a deeper understanding of LLMs [119]. At present, techniques such as chain-of-thought and program-of-thought can be employed to reveal the step-by-step reasoning and operational processes behind LLM outputs, improving their interpretability in medical applications [29, 120]. Researchers have also tried to dive into the inner workings of AI systems by introducing mechanistic interpretability, which is promising to provide a more interpretable and controlled LLM behavior [121].
The development, deployment, and production of LLMs in medical practice must adhere to strict regulatory standards and compliance requirements to ensure patient safety and data privacy [122]. Regulatory bodies such as the FDA, EMA, and others have established guidelines for medical devices, which increasingly include AI-driven tools [123]. Developers of medical AI applications must navigate these regulations, establish robust practices, and process procedures to ensure that their models are validated, transparent, and compliant with relevant laws, such as HIPAA in the U.S. Additionally, as AI applications are continuously updated and retrained, maintaining compliance over time requires ongoing monitoring, documentation, and potentially re-certification.
4.3 Evaluation and continuous improvement
@@
A medical AI system may comprise various modules and pipelines, each of which has the potential to malfunction, posing challenges in output assessment and debugging. Human oversight can be integrated to enhance validation, allowing users to confirm the accuracy of outputs through the provided references [124]. Furthermore, validation processes can be automated by leveraging LLM capabilities. For example, assertions can be embedded within the pipeline, enabling LLMs to self-correct their outputs during inference [125]. The increasing complexity of LLM-based systems, consisting of multiple interconnected components, can overwhelm any individual’s capacity to manage the entire architecture. LangSmith [126] is a DevOps tool for AI applications that aims to support deploying and monitoring LLM-powered applications. It allows developers to visualize traces and debug problematic components. Additionally, it facilitates the collection and extraction of erroneous inputs and outputs, which can then be utilized to build and augment validation datasets to enhance the applications.
Maintaining medical AI applications poses significant challenges in ensuring system stability and reliability. One major concern is the variability of the underlying LLMs, as version updates can alter model behavior and capabilities, potentially disrupting the functionality of integrated systems [127]. Additionally, the interdependence of system components means that changes to one element can impact the overall performance, necessitating rigorous testing and calibration whenever modifications are made. Furthermore, managing distribution shifts, particularly in edge cases, remains a critical issue. An effective strategy to address these challenges is the construction of robust development datasets for evaluating and fine-tuning AI pipelines. These datasets can be sourced from real-world tasks or simulated scenarios generated by LLMs [128]. However, current approaches often rely on heuristic methods, underscoring the need for further research to enhance dataset creation. Another significant challenge lies in selecting appropriate evaluation metrics, as many standards are embedded in specialized medical guidelines or regulatory documents. Bridging the knowledge gap between AI scientists and medical practitioners through increased collaboration and interdisciplinary research is essential for establishing reliable evaluation frameworks.
@@ 5 Conclusion
Generalist AI models have the potential to revolutionize the medical field. A range of adaptation methods have been proposed to tailor these models for specialized applications. In this Perspective, we documented existing adaptation strategies and organized them within a framework designed to optimize the performance of medical AI applications from a systems engineering perspective. Our analysis and discussion of published use cases demonstrate the benefits of this framework as a systematic approach to developing and optimizing LLM-based medical AI. However, we also recognize the potential challenges that arise as the complexity of medical AI applications increases, particularly in monitoring, validation, and maintenance. Future research and development are essential to solidify the utility of LLM-driven medical AI applications, enhance patient outcomes, democratize access to quality healthcare, and reduce the workload on medical professionals.